If you know an object's average angular speed and it's acceleration for an interval, can you determine the final angular velocity for that interval? Explain.
Answer #1

here

angular speed = 2 * pie / t

by using the first equation of the formula

v = u + a*t

angluar velocity in term of linear velocity is

w = v / r